OpenStudy (anonymous):

Let A be a diagonalizable matrix with every eigenvalue of A either 0 or 1. Prove that A is idempotent (A^2=A).

OpenStudy (anonymous):

This is what I've got so far: \[\text{By the definition of an eigenvalue, there exists nonzero vector } \vec x \text{ such that} \]\[A\vec x =\lambda \vec x \]\[\text{But }\lambda =\lambda^2 \text{ for all }\lambda=0,1. \text{ Hence, }\]\[A\vec x =\lambda^2 \vec x\]\[A\vec x=\lambda \left( \lambda \vec x \right)\]

OpenStudy (anonymous):

\[\text{Multiplying both sides by A on the \left}\]\[AA\vec x = A\lambda^2 \vec x\]\[\text{But }\lambda^2 = \lambda \text{ so we have}\]\[A^2 \vec x=A\lambda \vec x\]

OpenStudy (anonymous):

\[A^2 \vec x=\lambda A \vec x\]\[\text{But }A \vec x=\lambda \vec x \text{ so we have}\]\[A^2 \vec x=\lambda (\lambda \vec x)\]\[A^2 \vec x=\lambda^2 \vec x\]\[A^2 \vec x=\lambda \vec x\] Hence A is nilpotent

OpenStudy (anonymous):

Does that suffice?

OpenStudy (anonymous):

Correction: idempotent

OpenStudy (anonymous):

D with eigenvalues on the main diagonal yeah

OpenStudy (anonymous):

\[A^2=P^{-1}DP\]

OpenStudy (anonymous):

er D^2

OpenStudy (anonymous):

yeah that's much easier... I have no experience doing a lot of proofs

OpenStudy (anonymous):

hate going about it the long way when there's a simple solution

OpenStudy (anonymous):

thanks for your help

OpenStudy (loser66):

your prof asks you to prove it?

OpenStudy (anonymous):

yep

OpenStudy (loser66):

Prove theorem, it means you are junior math major in university?

OpenStudy (anonymous):

i'll minor in math